# Degree-Weighted DeGroot Learning

### v2

@@ 4.1 Influences and Wisdom of the Crowd
Example 1 (Influences when $\alpha=-1$).
When $\alpha=-1$, an agent $i$’s influence is given by
$s_{i}=\frac{\sum_{j}\frac{A_{ij}}{d_{i}d_{j}}}{\sum_{i^{\prime},j}\frac{A_{i^{\prime}j}}{d_{i^{\prime}}d_{j}}},$ (9)
which is inversely proportional to the (weighted) harmonic average of the neighbors’ degrees. When the neighbors’ degrees are independent, all agents share the same influence in expectation.
@@
Next, we highlight that the degree dependence, $\alpha$, via shaping the agents’ influences, can have more fundamental impacts on learning outcomes, such as determining whether the social learning process results in an unbiased aggregation of individual signals (i.e., the wisdom of the crowd).
Degree-Dependence and Wisdom.
@@
We follow the definition of wisdom introduced in Golub and Jackson (2010). More specifically, suppose that agents’ initial beliefs are i.i.d. signals drawn from a distribution with mean $\theta$, which represents the true state of the world, and variance $\sigma^{2}>0$. To study networks in the large, we index the learning matrix by $n$, representing the size of the society. The sequence of learning matrices is then given by a sequence of $n$-by-$n$ interaction matrices. We denote the belief of agent $i$ in network $n$ at time $t$ by $b_{i}^{(t)}(n)$ and its limiting belief by $b_{i}^{(\infty)}(n)$. We say that a sequence $\bigl{(}T(n)\bigr{)}_{n=1}^{\infty}$ of learning matrices is wise if the limiting beliefs converge jointly in probability to the true state $\theta$.
Definition 5 (Wisdom).
@@
The sequence $\bigl{(}T(n)\bigr{)}_{n=1}^{\infty}$ is wise if
$\lim_{n\to\infty}\max_{1\leq i\leq n}\Bigl{|}b_{i}^{(\infty)}(n)-\theta\Bigr{|}=0.$
@@
Let $s(n)$ denote the influence vector corresponding to $T(n)$, as defined in Proposition 1. By applying Lemma 1 and Corollary 1 in Golub and Jackson (2010) (with our formulation for $s(n)$ replacing their $s(n)_{1}$), we obtain an explicit necessary and sufficient condition for a society to be wise when agents use a degree‐dependent updating rule.
Corollary 1 (Corollary 1 of Golub and Jackson (2010)).
A society is wise if and only if
$\lim_{n\to\infty}\max_{1\leq i\leq n}s_{i}(n)=0.$
Below, we present two examples that illustrate how a shift in the degree‐dependent updating parameter $\alpha$ can transform a society from unwise to wise, and vice versa; we then generalize the insights derived from these examples. All proofs are relegated to Appendix A.
Example 2 (Unwise to Wise).
Consider a sequence of stochastic block models with two types of agents: one high-degree agent and $n-1$ low-degree agents. The linking probabilities are given by $p_{11}=0$, $p_{12}=p_{21}=1$, and $p_{22}=\frac{1}{n-1}$, where $p_{ij}$ denotes the probability that an agent of type $i$ connects with an agent of type $j$. The resulting expected degrees are $d_{1}=n-1$ for the high-degree agent and $d_{2}=2$ for the low-degree agents. Let $s_{1}$ denote the influence of the high-degree agent.
Then, a shift from $\alpha=0$ to $-1$ changes the learning outcome from unwise to wise. As $n\to\infty$:
  • For $\alpha=0$, $s_{1}\to 1/3$, non-vanishing; the learning is unwise.
  • For $\alpha=-1$, $s_{1}\to 0$; the learning is wise. [^20] The influence of each low-degree agent is always negligible, as there are infinitely many of them.
Example 3 (Wise to Unwise).
Consider a revised version of the previous example with the following linking probabilities: $p_{11}=0$, $p_{12}=p_{21}=\frac{1}{\sqrt{n-1}}$, $p_{22}=\frac{1}{n-1}$. The resulting expected degrees are $d_{1}=\sqrt{n-1}$ for the high-degree agent and $d_{2}=1+\frac{1}{\sqrt{n-1}}$ for the low-degree agents.
Then, a shift from $\alpha=0$ to $2$ changes the learning outcome from wise to unwise. As $n\to\infty$:
  • For $\alpha=0$, $s_{1}\to 0$; the learning is wise.
  • For $\alpha=2$, $s_{1}\to\frac{1}{2}$; the learning is unwise.
The observation in Examples 2 and 3—that a sufficiently strong change in degree dependence $\alpha$ can shift a society from wise to unwise (and vice versa)—extends to more general settings. As long as the high-degree agents constitute a bounded group with unbounded expected degree, and the low-degree agents form an unbounded group with bounded expected degree, the effect of $\alpha$ on societal wisdom is monotonic. We refer to such networks as high-degree–low-degree networks and summarize the general impact in the following proposition. [^21] See Appendix A for a more detailed description of the general setting.
Proposition 2 (Monotonic Impact of $\alpha$ on Wisdom).
Given any high-degree–low-degree network as described above, there exists a threshold $\bar{\alpha}>-1$ such that the learning outcome is wise if and only if $\alpha<\bar{\alpha}$. As a special case, the learning outcome is always wise when $\alpha\leq-1$.
Proof.
See Appendix A. ∎
Thus, degree dependence $\alpha$ has a monotonic effect on societal wisdom: higher values of $\alpha$ reduce the likelihood that society is wise. This monotonicity arises because wisdom obtains if and only if every individual’s influence vanishes. Non-wisdom emerges when an increase in $\alpha$ grants a certain group non-vanishing influence. In our setting, a sufficiently large $\alpha$ may allow a bounded group—such as high-degree individuals (e.g., experts or media)—to retain influence, rendering the society unwise. The result reflects a structural asymmetry: high-degree agents form a bounded group with unbounded links, while low-degree agents form an unbounded group with bounded links, ensuring that only the former can sustain influence in large societies.
Key Takeaway on Societal Wisdom.
@@
The above analysis shows that degree dependence $\alpha$ can significantly influence whether a society is wise or unwise. In networks with a bounded number of high-degree agents and infinitely many low-degree agents, the effect of $\alpha$ is monotonic: higher values of $\alpha$ grant high-degree agents non-vanishing influence and render the society unwise.
4.2 Degree-Dependence and Convergence Speed
@@
In this section, we illustrate the effects of the degree-weighted learning heuristic, $\alpha$, on convergence speed. Our key finding is that in the high-degree–low-degree-group network that we examine, the impact of $\alpha$ on speed is non-monotonic. Whether a more popularity-favoring rule (larger $\alpha$) accelerates or slows down convergence depends on the diversity of views potentially held by each group.
More specifically, we study a model with two types of groups: high-degree and low-degree. There may be multiple groups of each type, with high-degree groups having a larger expected degree than low-degree groups. [^22] See Appendix C for a detailed description of the setting.
Convergence speed, measured by $|\lambda_{2}|$, can be fully characterized in the high-degree–low-degree model. [^23] See Proposition 6 in Appendix C. We show that the effect of $\alpha$ on $|\lambda_{2}|$ is highly non-monotonic. Moreover, the specific impact of $\alpha$ depends on the number of high- and low-degree groups, $m_{h}$ and $m_{l}$.
Example 4 (One High-Degree Group and One Low-Degree Group).
Let $m_{h}=m_{l}=1$. Then, there exists a threshold $\alpha^{*}$ such that convergence is slowest at $\alpha^{*}$. As $\alpha$ deviates from this threshold, $|\lambda_{2}|$ decreases, implying faster convergence. [^24] See Theorem 2 in Appendix C for the characterization of $\alpha^{*}$.
@@
As illustrated in Example 4 (see the solid curve in Figure 2), changes in $\alpha$ lead to non-monotonic variations in learning speed. Convergence speed is not determined by a simple bias toward one group over the other; rather, the key determinant is the relative influence of the two groups. With a single high- and low-degree group, at most two dominant views emerge. When $\alpha$ is sufficiently large, the high-degree group dominates, leading to rapid convergence; likewise, as $\alpha\to-\infty$, the low-degree group dominates, also resulting in fast convergence. The slowest convergence occurs around an intermediate value $\alpha^{*}$, where the two groups exert relatively balanced influence.
Figure 2: Convergence Speed: One or Two High-Degree Groups and One Low-Degree Group
Notes: Convergence speed, measured by $|\lambda_{2}(\mathbf{T}^{*})|$, in degree dependence $\alpha$, for the following two cases:
One high-degree and one low-degree group (red solid): $n_{1}=10$ high-degree agents and $n_{2}=1000$ low-degree agents. The linking probabilities are $p_{11}=p_{22}=0.01$ within groups and $p_{12}=p_{21}=0.10$ across groups.
Two high-degree groups and one low-degree group (cyan dotted): $n_{1}=n_{2}=5$ high-degree agents and $n_{3}=1000$ low-degree agents. $p_{kk}=0.01$ within each group, $p_{k3}=p_{3k}=0.10$ for $k=1,2$ (between each high-degree group and the low-degree group), and $p_{12}=p_{21}=0.001$ between the two high-degree groups.
Example 5 (Two High-Degree Groups and One Low-Degree Group).
Let $m_{h}=2$ and $m_{l}=1$. The fastest convergence occurs as $\alpha\to-\infty$. For large positive values of $\alpha$, the pattern differs from Example 4, where there is only one high-degree group: convergence speed initially increases but then slows down as $\alpha\to+\infty$.
@@
Examples 4 and 5 (see the dotted curve in Figure 2) exhibit markedly different patterns, despite having nearly identical degree distributions. [^25] The parameters are chosen so that in both Examples 4 and 5, there are 10 high-degree agents with $d_{h}\approx 100$ and 1000 low-degree agents with $d_{l}=11$. With multiple high-degree groups, diverse views may arise among them. Unlike the single high-degree group case in Example 4, the limit $\alpha\to+\infty$ does not lead to immediate convergence, as conflicting high-degree groups require time to reconcile their differing views. [^26] For general cases with multiple high- or low-degree groups and explicit characterizations of $|\lambda_{2}|$ and the threshold $\alpha^{*}$, see Proposition 6 and Theorem 2 in Appendix C.
Key Takeaway on Convergence Speed.
@@
4.3 Trade-off Between Convergence Speed and Wisdom
Our analysis of degree-weighted learning’s impact on societal wisdom and convergence speed reveals potential trade-offs between the two, offering new economic insights. For instance, while a society that resists popularity bias may achieve greater wisdom, reaching consensus can take longer when many low-degree groups exist, each potentially holding different views. Conversely, a more popularity-favoring learning rule may accelerate belief convergence at the cost of wisdom.
We summarize the effects of $\alpha$ on wisdom and convergence speed across different cases in the following table:
Table 1: Trade-off between Wisdom and Speed under Different Popularity Conditions
# Groups Degree Dependence
High Low Popularity-resistant ($\alpha\to-\infty$) Popularity-favoring ($\alpha\to\infty$)
1 1 fast, wise fast, unwise
multiple 1 slow, wise slow, unwise
1 multiple slow, wise fast, unwise

### v3

@@ 4.1 Influence and Wisdom of the Crowd
Example 1
When $\alpha=-1$, an agent $i$’s influence is $s_{i}={\sum_{j}({A_{ij}}/{d_{i}d_{j}}})/{\sum_{i^{\prime},j}({A_{i^{\prime}j}}/{d_{i^{\prime}}d_{j}}}),$ which is inversely proportional to the (weighted) harmonic average of the neighbors’ degrees. When the neighbors’ degrees are independent, all agents share the same influence in expectation.
@@
Next, we highlight that the degree dependence, $\alpha$, via shaping the agents’ influences, can have more fundamental impacts on learning outcomes, such as whether the social learning process results in an unbiased aggregation of individual signals (i.e., the wisdom of the crowd).
Wisdom
@@
We follow the definition of wisdom introduced in Golub and Jackson (2010). More specifically, suppose that agents’ initial beliefs are i.i.d. signals drawn from a distribution with mean $\theta$, which represents the true state of the world, and variance $\sigma^{2}>0$. To study networks in the large, we index the learning matrix by $n$, representing the size of the society. The sequence of learning matrices is then given by a sequence of $n$-by-$n$ interaction matrices. We denote the belief of agent $i$ in network $n$ at time $t$ by $b_{i}^{(t)}(n)$ and its limiting belief by $b_{i}^{(\infty)}(n)$. We say that a sequence $\bigl(\mathbf{T}(n)\bigr)_{n=1}^{\infty}$ of learning matrices is wise if the limiting beliefs converge jointly in probability to the true state $\theta$.
Definition 4 (Wisdom)
@@
The sequence $\bigl(\mathbf{T}(n)\bigr)_{n=1}^{\infty}$ is wise if $\operatorname*{plim}\limits_{n\rightarrow\infty}\;\max_{i\leq n}\bigl|b_{i}^{(\infty)}(n)-\theta\bigr|=0.$
@@
We also say that a society is wise if the sequence $\bigl(\mathbf{T}(n)\bigr)_{n=1}^{\infty}$ is wise. Let $s(n)$ denote the influence vector corresponding to $\mathbf{T}(n)$, as defined in Proposition 1. By applying Lemma 1 and Corollary 1 in Golub and Jackson (2010) (with our formulation for $s(n)$ replacing their $s(n)_{1}$), we obtain an explicit necessary and sufficient condition for a society to be wise when agents use a degree‐dependent updating rule.
Lemma 1 (Golub and Jackson, 2010)
A society is wise if and only if $\lim\limits_{n\rightarrow\infty}\max_{i}s_{i}(n)=0.$
With the above preparations, we are now ready to state our results on the effect of degree dependence, $\alpha$, on influence and societal wisdom. Recall that for any expected network $\mathbf{R}$ generated by a two-type model introduced at the beginning of this section, there are two degree levels, denoted by $d_{l}<d_{h}$. Let the corresponding counts of agents be $n_{h}$ and $n_{l}=n-n_{h}$, and the influences of each type be $s_{l}$ and $s_{h}$, respectively. We have the following:
Proposition 2 (Monotonic Impact of $\alpha$ on Influences)
In the two-type model, as $\alpha$ increases, $s_{h}(\alpha)$ increases and $s_{l}(\alpha)$ decreases. In addition, at the extremes, we have $s_{h}(-\infty):=\lim_{\alpha\rightarrow-\infty}s_{h}(\alpha)=0$, $s_{l}(-\infty)=1/n_{l}$, and $s_{h}(+\infty)=1/n_{h}$, $s_{l}(+\infty)=0$.
That is, assigning more weight to high-degree neighbors amplifies the influence of high-degree agents and diminishes the influence of low-degree agents. At the extremes, when $\alpha\rightarrow$-$\infty$, high-degree neighbors are effectively ignored so that each low-degree agent carries equal influence, while the influence of any high-degree agent vanishes. Conversely, when $\alpha\rightarrow+\infty$, weight is placed almost exclusively on high-degree neighbors, so that each high-degree agent carries equal influence, while the influence of any low-degree agent vanishes.
With Lemma 1 and Proposition 2, we now state the comparative statics on wisdom.
Proposition 3 (Monotonic Impact of $\alpha$ on Wisdom)
In the two-type model, consider a sequence of societies indexed by the number of agents $n$, such that the number of high-degree agents is bounded, i.e., $n_{h}(n)<M$ for some $M<+\infty$. Then there exists a threshold $\bar{\alpha}$ such that the learning outcome is wise if and only if $\alpha\leq\bar{\alpha}$. At the extremes, the learning outcome is wise when $\alpha=$-$\infty$ and unwise when $\alpha=+\infty$.
Key Takeaway on Societal Wisdom
@@
In networks with a bounded number of high-degree agents and infinitely many low-degree agents, degree dependence $\alpha$ has monotonic effects on influence and thus on societal wisdom: higher values of $\alpha$ enable high-degree agents to retain non-vanishing influence, rendering the society unwise.
Wisdom requires that every individual’s influence vanish in the limit, whereas non-wisdom arises when a change in $\alpha$ grants a certain group non-vanishing influence. In our setting, as $\alpha$ increases, high-degree agents—who constitute a bounded group such as experts or media—may preserve influence, thereby rendering the society unwise. This reflects a structural asymmetry: while the group of low-degree agents grows unboundedly and their influence dissipates, the group of high-degree agents remains bounded and thus can sustain non-trivial influence in large societies (see Proposition 2).
4.2 Convergence Speed
In this section, we illustrate the effect of the degree-dependence $\alpha$ on convergence speed. The analysis builds on Proposition 1, which establishes that convergence speed is inversely related to $|\lambda_{2}|$ of the expected network, together with Theorem 1, which ensures that the comparative statics of $\alpha$ on convergence speed hold with high probability in large networks.
@@ 4.2 Convergence Speed
We show that, in contrast to the monotonic effect of $\alpha$ on wisdom, its impact on convergence speed is non-monotonic. Whether a more popularity-favoring rule (higher $\alpha$) accelerates or slows convergence depends on the relative number of groups, which essentially determines the diversity of views potentially held within each type.
Figure 2: Convergence Speed
{figurenotes}
Convergence speed, as measured by $|\lambda_{2}(\mathbf{T}^{*})|$, is plotted against degree dependence $\alpha$. The solid curve corresponds to the case with one high-degree group and one low-degree group (Example 2), while the dashed curve corresponds to the case with two high-degree groups and one low-degree group (Example 3).
We begin with the following example.
Example 2 (One High-Degree Group and One Low-Degree Group)
Consider a network consisting of one low-degree group (of size $n_{1}=1000$) and one high-degree group (of size $n_{2}=10$), where the within-group linking probabilities are $p_{11}=p_{22}=0.01$, and the across-group linking probabilities are $p_{12}=p_{21}=0.10$. There exists a threshold $\alpha^{*}$ at which convergence is slowest. As $\alpha$ moves away from this threshold, $|\lambda_{2}|$ of the expected learning matrix decreases, implying faster convergence.
@@
As the preceding example illustrates (see the solid curve in Figure 2), variations in $\alpha$ lead to non-monotonic changes in convergence speed. Notably, convergence speed is not simply governed by a uniform bias toward either group; rather, it is determined by the relative influence between the high- and low-degree groups. In the setting with one high-degree group and one low-degree group, at most two dominant opinions emerge. When $\alpha$ is sufficiently large, the high-degree group dominates, resulting in rapid convergence; similarly, as $\alpha\rightarrow$-$\infty$, the low-degree group prevails, again yielding fast convergence. The slowest convergence arises for an intermediate value $\alpha^{*}$, where the influence of the two groups is most balanced.
The following proposition generalizes Example 2 and provides an explicit characterization of the convergence speed for networks with two groups—one high-degree and one low-degree.
Proposition 4 (Convergence Speed with One High-Degree and One Low-Degree Group)
Consider a two-group model with $n_{1}>n_{2}$ and linking probabilities $p_{11}=p_{22}=p$ and $p_{12}=p_{21}=q$. The expected degrees of groups 1 and 2 are $d^{*}_{1}=n_{1}p+n_{2}q$ and $d^{*}_{2}=n_{1}q+n_{2}p$, respectively. The convergence speed can then be approximated by the inverse of the following expression:
(7) $|\lambda_{2}(\mathbf{T}^{*})|=\left|\frac{n_{1}p\,d_{1}^{*\alpha}}{n_{1}p\,d_{1}^{*\alpha}+n_{2}q\,d_{2}^{*\alpha}}-\frac{n_{1}q\,d_{1}^{*\alpha}}{n_{1}q\,d_{1}^{*\alpha}+n_{2}p\,d_{2}^{*\alpha}}\right|.$
This expression has the following properties:
  • $\lim\limits_{\alpha\rightarrow-\infty}|\lambda_{2}(\mathbf{T}^{*})|=0$, $\lim\limits_{\alpha\rightarrow+\infty}|\lambda_{2}(\mathbf{T}^{*})|=0$.
  • $|\lambda_{2}(\mathbf{T}^{*})|$ attains its maximum, $\overline{|\lambda_{2}|}={|p-q|}/{(p+q)}$, at $\alpha^{*}=\log_{d^{*}_{2}/d^{*}_{1}}(n_{1}/n_{2})$.
  • $|\lambda_{2}(\mathbf{T}^{*})|$ increases with $\alpha$ for $\alpha<\alpha^{*}$ and decreases with $\alpha$ for $\alpha>\alpha^{*}$.
The proposition reflects a general pattern of non-monotonicity in the impact of $\alpha$, as illustrated in Example 2, and shows that convergence is immediate when $\alpha\rightarrow+\infty$ or $\alpha\rightarrow$-$\infty$, as a single group of homogeneous agents (in expectation) dominates the learning dynamics.
The explicit characterization of $|\lambda_{2}|$ also provides new insight beyond the example: as shown in the proposition, the slowest rate of convergence is captured by $\overline{|\lambda_{2}|}={|p-q|}/{(p+q)}$. Recall that $p$ denotes the linking probability within groups and $q$ the linking probability across groups. When $p>q$, this recovers the classical result that increased homophily slows convergence (Golub and Jackson, 2012). More interestingly, when $p<q$, the slowest rate of learning decreases in |pq||p-q|, indicating that greater group differentiation also slows convergence. Finally, if $p=q$, the groups are indistinguishable and the population is homogeneous (in expectation), yielding immediate convergence for any value of $\alpha$.
To further illustrate that it is the diversity of views that drives the non-monotonic impact of $\alpha$, we consider the following example, which splits the single high-degree group in Example 2 into two groups.
Example 3 (Two High-Degree Groups and One Low-Degree Group)
Let $n_{1}=1000$ and $n_{2}=n_{3}=5$, with linking probabilities $p_{kk}=0.01$ within each group, $p_{1k}=p_{k1}=0.10$ for $k=2,3$ between the high-degree and low-degree groups, and $p_{23}=p_{32}=0.001$ between the two high-degree groups. The learning speed for this configuration is plotted as the dashed curve in Figure 2. For small values of $\alpha$, it closely resembles Example 2 (the solid curve). For large positive $\alpha$, however, the pattern diverges markedly: convergence slows down as $\alpha\rightarrow+\infty$.
@@
Examples 2 and 3 feature almost identical degree distributions, [^27] The parameters are chosen so that in both examples, there are 10 high-degree agents with expected degree $d_{h}\approx 100$ and 1000 low-degree agents with expected degree $d_{l}\approx 11$. yet exhibit markedly different patterns of convergence speed. Unlike the single high-degree group case (Example 2), with multiple high-degree groups (Example 3), diverse views may arise among them; consequently, as $\alpha\rightarrow+\infty$, immediate convergence does not occur, since conflicting high-degree groups require time to reconcile their differing opinions. The following proposition generalizes Example 3 to cases with multiple high-degree or multiple low-degree groups.
Proposition 5 (Convergence Speed with Three or More Groups)
The learning speed, captured by $1/|\lambda_{2}(\mathbf{T}^{*})|$, is generally non-monotonic in the learning heuristic $\alpha$. [^28] For explicit characterizations of $|\lambda_{2}|$ and the threshold(s) of $\alpha$, see Lemma C.1 and Proposition C.1 in Appendix 2. In particular:
  • $m_{h}>1,\,m_{l}=1$: There exist two thresholds, $\alpha_{1}<\alpha_{2}$, such that $|\lambda_{2}(\mathbf{T}^{*})|$ increases with $\alpha$ when $\alpha<\alpha_{1}$ or $\alpha>\alpha_{2}$, and decreases with $\alpha$ when $\alpha\in(\alpha_{1},\alpha_{2})$. In addition, $\lim\limits_{\alpha\rightarrow-\infty}|\lambda_{2}(\mathbf{T}^{*})|=0$, $\lim\limits_{\alpha\rightarrow+\infty}|\lambda_{2}(\mathbf{T}^{*})|>0$.
  • $m_{h}=1,\,m_{l}>1$: There exist two thresholds, $\alpha_{1}<\alpha_{2}$, such that $|\lambda_{2}(\mathbf{T}^{*})|$ decreases with $\alpha$ when $\alpha<\alpha_{1}$ or $\alpha>\alpha_{2}$, and increases with $\alpha$ when $\alpha\in(\alpha_{1},\alpha_{2})$. In addition, $\lim\limits_{\alpha\rightarrow-\infty}|\lambda_{2}(\mathbf{T}^{*})|>0$, $\lim\limits_{\alpha\rightarrow+\infty}|\lambda_{2}(\mathbf{T}^{*})|=0$.
Key Takeaway on Convergence Speed
@@
4.3 Trade-off Between Wisdom and Convergence Speed
Our analysis of the impact of degree-weighted learning on societal wisdom and convergence speed reveals a potential trade-off between the two. For instance, while a society that resists popularity bias may achieve greater wisdom (see Proposition 3), consensus may be reached more slowly when many low-degree groups exist, each potentially holding different views (see the second case in Proposition 5). Conversely, a more popularity-favoring learning rule may accelerate belief convergence at the cost of wisdom.
Table 1, introduced in the Introduction, summarizes the results on societal wisdom (wise or unwise, Proposition 3) and convergence speed (fast or slow, Propositions 4 and 5). We discuss potential policy implications of the wisdom–speed trade-off in the concluding remarks.